News Alert: Situational Awareness Report 8 February 2017 As of 3:00 pm
Posted: 2/8/2017 3:11:37 PM
Louisiana Business Emergency Operations Center

Situational Awareness Report

8 February 2017

As of 3:00 pm

 

State EOC Activation Level: Level III

WebEOC Incident Name: 17-008 Severe Wx — Statewide — Feb07

State Declaration Proclamation 17 JBE 2017 State of Emergency – Severe Weather

 

 

Situation: Multiple tornadoes spawned from the severe weather that moved across the state yesterday. Heavily affected areas include parts of Livingston, St. James, Ascension, Tangipahoa, St. Tammany, Jefferson, and Orleans Parishes, including eastern New Orleans. Over the next few days, the National Weather Service will be assessing the amount, intensity, and paths of tornadoes in this outbreak.

 

 

***All updates to this report will be in red text***

 

STATE AGENCIES

 

GOHSEP

·        GOHSEP is activated to a Level III

·        State damage assessment teams have begun preliminary damage assessments this morning (Wednesday, February 8).

·        A WebEOC Incident has been created for this event.

·        Regional Coordinators are maintaining communication with parishes and GOHSEP for situational awareness.

·        The Governor issued an emergency declaration (17 JBE 2017).

 

Louisiana National Guard

·        Will provide security at 25 static checkpoints throughout the damaged area, along with helicopter support by two Blackhawks.

 

Louisiana State Police

·        TROOP A STAFFING: A total of 16 Troopers, including 1 Lieutenant and 2 Sergeants.  A total of 4 Troopers were sent to Donaldsonville to assist the Ascension Parish Sheriff’s Office with tornado damage searches and recovery.

·        TROOP B / TROOP N STAFFING: A total of 35 Troopers are in Orleans Parish, including 1 Trooper at the Orleans Parish Office of Emergency Preparedness. The area directly impacted by the tornado is a residential area within the boundaries of Wilson Avenue, Dwyer Road, Wright Road, and Chef Menteur Highway in New Orleans East. Troopers are conducting searches of damaged houses, manning roadblocks, and assisting with looting patrols in the area.

·        TROOP L STAFFING: Troop L has 2 Sergeants, and 9 Troopers assigned to the current shift. Two of those Troopers have been sent to New Orleans East to assist with storm damage operations. Team E has 1 Lieutenant, 1 Sergeant and 1 Trooper currently working. Communications is being handled out of HQ as of 2pm. Troopers are actively monitoring roadways and are in contact with local agencies.

·        CRISIS RESPONSE: Crisis Response Staff providing SITREPS to the GOHSEP SEOC and monitoring WebEOC for requests.  The LSP EOC is not activated at this time.

 

State Fire Marshal

·        2/7/17 OSFM Personnel from N.O. Area 1 continue to support ongoing operations with NOFD. NOFD assembled 8 USAR search teams with a medical specialist assigned to each group for a total of 40 members at 1600 hours to provide a secondary search of the affected area. The search area is between Wilson to Bullard and Chef to Dwyer.  Search teams were briefed and assigned zones at 1630 hours. The teams deployed at 1700.  By 1845, the teams were reporting approximately 25% of the secondary search complete. The teams will re-evaluate the progress at 1930 and the operation' chief will determine if the secondary search will continue.

·        As of 0600 2/8/17 NOFD is back and operational at CP on Wilson and Chef Menteur Roll call: Melanson, Santos, Patel, Correa, Davenport, Verges

·        OSFM 4 man team has just finished a recon of grid north of I-10 adjacent  to lakefront airport for unconfirmed reports of additional damage(Unfounded) OSFM assisted in a complaint inspection of a 6story mixed income/senior citizen apartment building with loss of power . Structure was outfitted with a diesel generator which failed after transfer from loss of shore power. NOEMS TAC medics and OSFM attempted to assess generator until an electrician could be located. Licensed electrician arrived at Owner request and determined the generator was beyond basic field repair. Due to the impractical feasibility to relocate ambulatory residents a fire watch was ordered by this office and remains in place by NOFD. Generator maybe repaired as early as this AM.

·        LANG&LSP continue to provide field security

·        No reports of injuries or transports of civilians or field personnel.

 

Civil Air Patrol

·        Conducting flights across southeast Louisiana to assist with damage assessment.

 

Public Service Commission

·        As of 10:00, there were 8,573 power outages remaining due to tornadoes/severe weather, with a parish breakdown of:

·        Ascension: 804

·        Orleans: 7,452

·        St. James: 38

·        St. Tammany: 279

·        New Orleans East: Entergy crews made tremendous progress overnight restoring power to approximately 4,000 customers.  Crews continue to assess the damages and work to restore power in the New Orleans East area stretching from Lake Forest through Sherwood Forest to Michaud along Chef Meteor Blvd.  Contractors have arrived in the area to assist Entergy with damage assessment and restoration efforts.

·        There are approximately 250 Entergy crew members and contractors working in the area to assess damage, clear debris in roads and repair the electrical facilities. Crews have been able to complete damage assessments to 70% of the area. Preliminary damage reports include roughly 160 damaged and downed poles, 100 damaged transformers, 50 downed transmission poles and 150 spans of downed and damaged distribution lines. While we will continue to restore customers throughout each day, the completion of the restoration is expected to take 3-5 days.

·        Schools without power in New Orleans East include Schaumburg Elementary and Einstein Charter School Sherwood Forest.

·        Entergy gas structures in New Orleans East were also damaged.  Approximately 40 Entergy gas crew members have assessed the damage and responded to 69 leak calls ensuring locations are safe. Walkthroughs were completed in the most heavily damaged neighborhoods. Approximately 200 customers have had their gas service turned off due to leaks or severe structural damage. Entergy will restore gas service as soon as customers are ready as long as their house fuel line piping is undamaged and can pass the safety check.

·        Damage assessment has been completed for 70% of Entergy’s facilities on the west bank of Ascension and St. James parishes.  Entergy has brought in 50 additional restoration workers to assist with the restoration in the area. Damage assessment identified over 45 broken poles in the region. For customers who can safely accept power, service restoration is anticipated for this evening, Wednesday, Feb 8.  Entergy Customer Service Representatives are working directly with DCFS already to get them lists of customers without power for 24 hours or more, and they are working on getting reps in the NOLA area to help do the same there.

·        Cleo’s system in the Port Louis area has been devastated by tornados.  Cleo has six broken concrete poles located in the marsh that must be repaired before power can be restored to 56 customers in the area. Cleo is securing specialized equipment to reach the damaged part of its system. Cleo anticipates power could remain out through Friday, Feb. 10.

 

Department of Children and Family Services

·        Shelter Population is 78 as of 2/8/17 1200. Faith Christian Center closed at 12noon today.

·        DCFS is working to provide information on the process to be used for current SNAP recipients who have lost food due to tornado damage or power outages.

 

Department of Health and Hospitals

·        High level ‘take-aways’:

o   Shelter was opened in Region 1 in the New Orleans East area as approximately 75 homes were damaged and displaced families (3 shelterees as of 3:49 pm)

o   28 injuries absorbed into Region 1 Hospitals (as of 3:49 pm)

o   9 injuries absorbed into Region 2 Hospitals

o   2 injuries absorbed into Region 9 Hospitals

o   Search and Rescue assessments are still underway – which may generate additional injuries.

·        General ESF8 Response Activities:

o   Alerts were sent out via ESF8 Portal. 

o   LERN directed trauma transports.

o   Reports came in from Louisiana Emergency Response Network (LERN), Hospital and EMS DRCs and Public Health Emergency Response Coordinators (PHERC) and were compiled by EOCWatch Desk.

o   For CMS purposes, Health Standards Section is reaching out to ASCs, RHC, FQHCs, CMHC, HH, hospice, PRTFs, ICFs, CORF, OPT, ESRD to determine impact. LNHA and Dialysis Network are also reaching out to their members to determine impact. As you may know, it is required for licensed facilities to contact HS if they are impacted by a disaster.    

 

American Red Cross

·        Shelter Open: Joe Brown Memorial Park Reception Center changed to Shelter.  Parish Managed.  Red Cross logistically supporting with food and supplies- 78 Population as of 8:00am

·        Faith Christian Center Shelter in Ascension Parish closed at 12:00 pm today.

·        8:00am Disaster Leadership Operation Call this morning held.  Disaster Leadership Call scheduled for this afternoon

·        IAPv1 distributed to State and Federal government partners and VOAD Lead for State

·        SEOC -  State EOC Activation Level: Crisis Action Team (CAT).  Supporting virtually and monitoring WebEOC

·        Red Cross communicating with affected local OEPs

·        Provided dinner for 50 people at the New Orleans EOC last night

·        Public Affairs continues to send out safety messages and updating news social and news media

·        Continue to feed Joe Brown Park Shelter and New Orleans Mobile Command Post

·        Meeting scheduled today with ARC Leadership and Orleans Parish Leadership to discuss shelter management transition

·        Mobile Canteening will continue today in the New Orleans East area

·        Red Cross DHS/DMH/Spiritual Care will be supporting community today

·        DA will continue today in the affected areas.  Will provide update this afternoon on our DA Assessment

·        Continue to support New Orleans EOC with ARC Liaison

 

Emergency Management Disability and Aging Coalition

·        EMDAC held an Access and Functional Needs Conference call today at 12:30

·        EMDAC members can assist with lost DME, Medical supplies, and help restore medical services.

·        Governor's office of Elderly Affairs has reached out to the Councils on Aging to see if any of their clients have any unmet needs.  There have been no reports of needs at this time.

·        EMDAC has reached out to New Orleans to offer services and help with anyone who may need assistance with lost DME or Medical Supplies.  EMDAC members can conduct shelter assessments if the parish needs them to do it.  ADA administrator for the City of New Orleans participated on the call.

·        ARC of Louisiana is reaching out to the local ARCs to see if they have any unmet needs.

 

Louisiana Department of Environmental Quality

·        Tornado went through the area and damaged storage tanks, buildings and power lines. 

·        A natural gas line was also damaged by tornado, but gas is off, not blowing. -  Material released from storage tanks into containment. In process of putting boom out around containment area as a precaution.  Power is currently out.

·        LDEQ Emergency Response still on site at Crystal Clean Facility

·        Above ground used oil tank leak slowed significantly -Wastewater tank still leaking -All used oil and wastewater released inside secondary containment with no off-site impact. Clean-up crews on site. No injuries, fire, or road closures due to incident

 

Department of Transportation and Development

·        DOTD ESF-3 Public Works and Engineering is in contact with DOTD Districts in impacted areas, DODT District personnel are coordinating with local OEPs and public works.

·        DOTD District 02 personnel has advised that they are working to clear US 90 in New Orleans East. They are with the Orleans command group in the field coordinating response. DOTD District 02 has been requested to provide 50 type III barricades and cones to assist with road closures.  DOTD District 02 has the devices in route and employees to install.  

·        Providing OSFM with two mobile light towers for search and rescue in the New Orleans area.

 

Department of Agriculture and Forestry

·        LDAF is monitoring WebEOC and parish animal control agencies for resource request. No animal issues reported at this time.   

 

Salvation Army

·        The Baton Rouge Corps is deploying a unit to Donaldsonville in Ascension Parish today.

·        New Orleans Area Command: One (1) Canteen (mobile kitchen) on scene in impacted neighborhood in East New Orleans served 112 meals, 95 snacks and 266 drinks.  It will remain on scene until further notice.  A second local unit is on standby.

·        Baton Rouge Corps: A Canteen (mobile kitchen) was deployed to Watson, LA, in Livingston Parish.  It served 75 meals.  Meal service will be continued on Wednesday, 08 February  

 

Louisiana VOAD

·        Louisiana VOAD is monitoring and on-standby to respond as requested.  Mercy Chefs mobile feeding is on standby and ready to respond as needed for emergency/community feeding.

·        Samaritan's Purse in route to New Orleans for assessment. 

·        Crisis Cleanup has a live incident "S LA Tornadoes Feb. 2017" available to coordinate work orders for cleanup.    

 

 

PARISH ACTIONS

 

Region 1

 

Orleans

·        Tornado activity was spotted by NOFD and NOPD in the area of Chef Highway and Downman Road. Parish field surveys are underway, however current indication is EF 2 or 3.

·        More than 50 homes with moderate to heavy damage near Chef Highway and Goode Street, More than 20 homes with moderate to heavy damage near 4600 block of Nottingham near Camelot.

·        Einstein Sherwood Forest School and Schaumburg School both sustained heavy damage. No injuries but children are sheltering in place and a dismissal program is being put together now.

·        At NASA MICHOUD sustained damage to final assembly side of facility, South side of building 103, and approximately 80 percent of cars in the parking lot flipped or damaged. The USDA building was hit hard, gas leaks being identified and shut down. Delaying securing flight hardware until we're assured the building structure is sound. MAF is currently under another shelter in place.

·        Approximately 5,698 are still without power.

·        Approximately 31 injuries.

·        Joe Brown Park Shelter: 93 pax have come through for services, 78 identified as planning on staying in the shelter

·        Debris and damage assessments began yesterday and have resumed as of 0600 this morning.

·        A LANG LNO for the city EOC has been established. LANG is also providing additional security assets. RTA service has been modified on Chef Manteur due to storm damage.

·        7 schools are closed today in New Orleans East due to power outages.

 

Jefferson

·        NWS confirmed tornado touchdown at approximately 10:45 am 2/7/17 between the 700 block of St George Avenue and 3700 block of Jefferson Highway. Parish will continue to work with NWS to confirm path and intensity. JPEM has been working closely with the NWS reporting areas of remnants from structural damage and visible flying debris from the Old Jefferson neighborhood.  A complete report will be forwarded for examining events by the NWS field response forecasters to determine the strength of  tornado. 

·        Approximately 20 structures sustained light to moderate damage.

·        Power outages are down to 250 Entergy customers.

·        All streets are cleared from debris and flood waters and have no travel issues at this time.

·        No injuries to report.

 

Region 2

 

Livingston

·        Officials and NWS confirmed tornado activity in the Killian and Springfield area.

·        5 homes have been damaged and 2 destroyed with numerous trees and powerlines down in the Killian area.

·        Watson area damages as confirmed by OSFM include 5 homes heavily devastated (3 of which were mobile homes), 1 camper trailer, and several vehicles.

·        3 injuries reported, 2 of which were ambulanced away and 1 stayed.

·        Emergency Medical Services and American Red Cross responded.

 

Ascension

·        Tornado confirmed to have touched down in the Donaldsonville area at approximately 11:00 am 2/7/17 near the Old Rice Mill and northeast into the city including Williams and St. Patrick Street. The Director assessed damages, so far those reveal roof damage, car damage and vegetative debris.

·        CF Industries have also reported some damages.

·        Approximately 1,700 without power.

·        3 injuries in St Gabriel transported to area hospitals.

 

Region 3

 

St. James

·        NWS started assessment at approx. 0930 hrs.

·        PDA teams enroute to affected areas for their assessment at 1200 hrs.  

·        Tornado reported west side of the Sunshine Bridge on June Street in a small community known as Lemannville. NWS is scheduled to conduct their analysis today.

·        The Advocate reported a confirmed tornado in the Lutcher area, OHSEP stating this is NOT TRUE.

·        13 homes damaged with an additional 5 destroyed.  Law Enforcement, Parish Social Services and EM staff conducted  assessments of residents and damaged areas. Crews are hindered with assessments due to downed trees and power lines.

·        95 % of Entergy customers have been restored. Power to customers in Lemannville area should be restored by this evening.

·        2 reported minor injuries.

·        American Red Cross responded. Sheriff’s Office is providing 24 hour security; Sheriff’s mobile command is located on Charles Street. Public Works crews will assist in debris removal. Volunteers are also assisting residents with minor roof repairs.

·        Gas and Water services are disconnected until parish utilities deem safe for restoration. Parish will go door-to-door to advise residents on restoring power.

·        Parish EOC returned to normal operations.

 

Region 9

 

Tangipahoa

·        OHSEP reported 3 possible tornadoes impacted the parish. First, a large tornado moved through Akers south of Ponchatoula heading towards Madisonville. A second that caused damage in Ponchatoula and possible a third near Amite. NWS believes this may be the same as what passed through the Killian area.

·        Parish reported 8 power poles down on Owl Bayou.

·        No injuries to report.

 

St Tammany

·        Reports of a tornado touchdown in Madisonville at Highway 22 and Pine Street with possible damage in northern St Tammany.

·        3 homes in the Madisonville area with moderate to major damage.

·        No injuries to report.

The LA BEOC Portal is designed and developed by the NIMSAT Institute under the Cooperative Endeavor Agreement with the Louisiana Governor's Office of Homeland Security and Emergency Preparedness (GOHSEP)

GOHSEP LED NIMSAT

© Copyright 2010-2018, NIMSAT. All rights reserved. Terms of Service. Portal Guidelines.